OpenStudy (anonymous):

please explain..Under water, pressure increases 4.3 pounds per square inch (psi) for every 10 feet you descend. This is represented by the equation p = 0.43d + 14.7, where p is the psi and d is the depth in ft. Find the pressure at 200 feet. 215.13 psi 860 psi 100.7 psi

OpenStudy (anonymous):

p = (0.43)(200) + 14.7

OpenStudy (anonymous):

Notice how this is a linear equation in the form p = md + b which is similar to what you already work with as y = mx + b

OpenStudy (anonymous):

So, "d" acts like "x" in that it is your independent variable and is used or identified first to get "p". So, "p" depends on "d".

OpenStudy (anonymous):

uuuuuh?

OpenStudy (anonymous):

can you show in steps?

OpenStudy (anonymous):

ok. Your initial equation is p = (0.43)d + 14.7 and you go down 200 feet. So 200 is your "d" so you substitute for that to get p = (0.43)(200) + 14.7 and then just multiply the (0.43)(200) to get 86 and then add 14.7

OpenStudy (anonymous):

And then 86 + 14.7 = 100.7 And that's all there is to it!

OpenStudy (anonymous):

Making sense?

OpenStudy (anonymous):

One last thing that might need explanation. Pressure increases 4.3 every 10 feet. That is the same ratio as 0.43 for every foot. That's why the equation has 0.43 instead of 4.3 in it.

OpenStudy (anonymous):

yuip thanks once again:)

OpenStudy (anonymous):

you're welcome!